OpenStudy (bananas1234):

Simplify. (a^3b^–4)^3

OpenStudy (bananas1234):

options: 1/a^3b^3 1/ab a^9/b^12 a^6/b^7

jagr2713 (jagr2713):

ok first we use distributive property (Multiplication) which is\[(xy)^{a}=x ^{a}y y^{a}\] can you set it up?

jagr2713 (jagr2713):

If not here it is\[(a ^{3})^{3}(b ^{-4})^{3}\] what o we do with the two 3's in the equation @Bananas1234

jagr2713 (jagr2713):

You the @Bananas1234

OpenStudy (anonymous):

Um.. idk the answer... :P

jagr2713 (jagr2713):

@Bananas1234 you there

OpenStudy (bananas1234):

yes, sorry i had to step away for a min.

OpenStudy (bananas1234):

add the 3s together to get a^6?

jagr2713 (jagr2713):

Nope we simplify right? what do we get? and give it to me in the equation form

OpenStudy (bananas1234):

we could divide them then all that would be left is a?

jagr2713 (jagr2713):

Nope simplify means or multiply :D

OpenStudy (bananas1234):

a^9?

jagr2713 (jagr2713):

Good so we are left with??

OpenStudy (bananas1234):

(b^-4)^3

jagr2713 (jagr2713):

you forgot the a9

jagr2713 (jagr2713):

so \[a ^{9}(b ^{-4})^{3}\] so what do we do with the next sie the -4 and 3

OpenStudy (bananas1234):

multiply or divide?

jagr2713 (jagr2713):

Its like we did above :D and put it into the equation like above

OpenStudy (bananas1234):

b^12?

jagr2713 (jagr2713):

Your missing something again (-4)(3)

OpenStudy (bananas1234):

-12?

jagr2713 (jagr2713):

Good so we have \[a ^{9}b ^{-12}\]

jagr2713 (jagr2713):

Ok now we use Negative Power Rule \[x ^{-a}=\frac{ 1 }{ x^{a} }\] Can you set it up as that?

OpenStudy (bananas1234):

a^-9 = 1/b^-12? sorry if that is wrong

jagr2713 (jagr2713):

Its ok is the 9 Negative?

OpenStudy (bananas1234):

no

jagr2713 (jagr2713):

So its actually?

OpenStudy (bananas1234):

b^12 = 1/a^9

jagr2713 (jagr2713):

Good so when we simply we get or just put it into a fraction what do we have?

OpenStudy (bananas1234):

a^9/b^12

jagr2713 (jagr2713):

Correct :D There you go :D

jagr2713 (jagr2713):

Great job :D

jagr2713 (jagr2713):

so whats the answer?

OpenStudy (bananas1234):

C

jagr2713 (jagr2713):

Correct

OpenStudy (bananas1234):

Thank you!

jagr2713 (jagr2713):

Anytime :D